OpenStudy (anonymous):

Use polar coordinates to set up and evaluate the double integral: e^-(x^2+y^2)/2

OpenStudy (anonymous):

\[f(x,y)=e^-(x^2+y^2), R: x^2+y^2\le25,x \ge0\]

OpenStudy (anonymous):

For the limits I got \[y=\sqrt{x^2-5}\], y = 0, x=5, x=0

OpenStudy (anonymous):

So, \[\int\limits_{0}^{5}\int\limits_{0}^{\sqrt{x^2-5}}e^{-(x^2+y^2)/2}dydx\]

OpenStudy (anonymous):

This is how far I've gotten right now, I think next I have to convert the integrals to polar coordinates.

OpenStudy (experimentx):

let x^+y^ = r^2 and dxdy = r dr d\(\theta\) It should be pretty easy now.

OpenStudy (anonymous):

alright, I'll try solving this and let you guys know what I get or if I get stuck somewhere.

OpenStudy (anonymous):

^ Since I'm converting the limits would change right? If so would they be \[d \theta=0,pi/2, r=0,5\]

OpenStudy (anonymous):

dθ=-pi/2,pi/2,r=0,5

OpenStudy (anonymous):

Oh okay...

OpenStudy (anonymous):

\[\int\limits_{-pi/2}^{pi/2}\int\limits_{0}^{5}r^2(\cos^2\theta+\sin^2\theta)r dr d \theta\] \[1/8\int\limits_{-pi/2}^{pi/2}-r ^{4} d \theta\] \[1/8\int\limits_{-pi/2}^{pi/2}625d \theta\] ^I got this by plugging in 5 and 0.

OpenStudy (anonymous):

That's what I get when I evaluated with respect to r. I have a feeling that I should of brought that - sign over with the 1/8

OpenStudy (anonymous):

\[e^{(x^2+y^2)/2}\] or \[e^{x^2+y^2}/2\] ? @JerJason

OpenStudy (anonymous):

The first one, sorry I forgot to change that when I originally posted it

OpenStudy (anonymous):

ah ok...  so you see: 0<=x<=5 -sqrt(25-x^2)<=y<=sqrt(25-x^2) so we'll have the double integral \[\int\limits_{0}^{5} \int\limits_{-\sqrt{25-x^2}}^{\sqrt{25-x^2}}e^{-(x^2+y^2)/2}dydx\] now looking at this with polar coordinates....  we can see that -pi/2<=theta<=pi/2 0<=r<=5 recall that x^2+y^2=r^2 so e^(-(x^2+y^2)/2)=e^(-r^2/2) so we'll have the double integral: \[\int\limits_{-pi/2}^{pi/2} \int\limits_{0}^{5}e^{-r^2/2}rdrd \theta\]